# Inflationary Gravitational Wave Spectral Shapes as test for Low-Scale Leptogenesis

### v1

@@ III.2 The Boltzmann Equations
The evolution of the energy densities of matter and radiation and the number density of $B-L$ asymmetry are determined by the Boltzmann equations,
$\dot{\rho}_{\phi}$ $=-3\mathcal{H}\rho_{\phi}-\Gamma_{\phi}\,\rho_{\phi}$ (30)
$\dot{\rho}_{N_{1}}$ $=-3\mathcal{H}\rho_{N_{1}}+\Gamma_{\phi\rightarrow N_{1}N_{1}}\,\rho_{\phi}-\Gamma_{N_{1}}(\rho_{N_{1}}-\rho_{N_{1}}^{eq})$
$\dot{\rho}_{N_{2}}$ $=-3\mathcal{H}\rho_{N_{2}}+\Gamma_{\phi\rightarrow N_{2}N_{2}}\,\rho_{\phi}-\Gamma_{N_{2}}(\rho_{N_{2}}-\rho_{N_{2}}^{eq})$
$\dot{n}_{B-L}$ $=-3\mathcal{H}n_{B-L}-\epsilon\sum_{i=1}^{2}\Gamma_{N_{i}}(n_{N_{i}}-n_{N_{i}}^{eq})-\Gamma_{ID}\,n_{B-L}$
$\dot{\rho}_{R}$ $=-4\mathcal{H}\rho_{R}+\Gamma_{\phi\rightarrow R}\,\rho_{\phi}+\sum_{i=1}^{2}\Gamma_{N_{i}}(\rho_{N_{i}}-\rho_{N_{i}}^{eq})$
where $\mathcal{H}$ is the Hubble parameter in cosmic time and dot denotes derivative with respect to cosmic time and $\rho^{eq}_{X}$ and $n_{X}^{eq}$ represents the equilibrium energy and number density of the species $X$ respectively. $\Gamma_{ID}$ is the inverse decay rate. See Appendix A for details. Here we have considered only decays and inverse decays and neglected scattering processes of the RHNs and the scalar. Also, for simplicity here we considered only non-relativistic RHN production in the equations for $N_{1,2}$, which may not be guaranteed when $M_{\phi}\gg M_{1,2}$.
While solving the Boltzmann equations we take the coupling $y_{R}=0$ defined in Eq. (19) such that $\Gamma_{\phi\rightarrow R}\sim 0$ in order to maximize the RHN production from $\phi$ and hence the lepton asymmetry. Later we will consider the $y_{R}>0$ scenario which is similar to thermal leptogenesis. We start solving the Boltzmann equations from the time when the scalar energy density overtakes the radiation energy density. Thus the initial energy density of the scalar is equal to the radiation density Kolb:1990vq ,
@@
We take $T_{\rm dom}$ from Eq. (25). The initial densities of RHNs are taken to be free parameters smaller than $\rho_{\phi}(T_{\rm dom})$ while initial $B-L$ asymmetry is taken to be negligible. Further details on the numerical solution of the Boltzmann equations and calculation of final $\kappa_{f}$ value as-well-as maximum available temperature are given in Appendix A. In Fig. 2 along with the thermal case we also show the enhancement in $\kappa_{f}$ in non-thermal leptogenesis from $\phi$ decay by solving the Boltzmann equations (30). The solid magenta line, dashed green line and the dash-dotted purple line shows the final efficiency factor as a function of the washout parameter for a special case, $M_{1}\sim T_{\phi}$ Hahn-Woernle:2008tsk . Here we took $M_{1}=10^{5}$ GeV, $M_{\phi}=10^{8}$ GeV and $T_{\phi}=10^{5},5\times 10^{4},10^{4}$ GeV. Efficiency factor is maximum at $K\sim 0.3$ for $M_{1}\sim T_{\phi}$. We note that for $T_{\phi}\ll M_{1}$, the final efficiency factor becomes independent of $K$.
At this point we are able to numerically calculate the final baryon asymmetry in our model. However we are not interested in the exact amout of asymmetry generated, since our goal is studying GW probe of leptogenesis. In the next section we discuss classification of scenarios based on numerical results and analytical approximations, relevant for GW sensitivity. We make sure that for all cases, observed BAU is successfully generated.
Figure 2: Analytical estimate of final efficiency factor $\kappa_{f}$ as a function of washout parameter $K$ in thermal and numerical solutions to the Boltzmann equations (30) in non-thermal leptogenesis. Red solid curve represents thermal initial abundance of $N_{1}$ and blue dashed curve represents vanishing initial $N_{1}$ abundance in thermal leptogenesis. Here we used Eq. (13) and (14) for the blue curve while for the red curve we use Eq. (12) for $K\ll 1$ and interpolate it to the thermal case in the $K\gg 1$ regime. The solid magenta, dashed green and dash-dotted purple lines represent non-thermal leptogenesis with $M_{1}=10^{5}$ GeV, $M_{\phi}=10^{8}$ GeV for different $T_{\phi}$ values.
@@ IV Classification of Scenarios
In this section, we identify two general non-thermal scenarios relevant for GW observations along the lines of Zhang:2023oyo . Additionally we introduce a thermal leptogenesis scenario with two-step entropy injection. We present analytical estimates of the equilibrium temperature when matter domination ends, and the dilution factor from entropy injection for each of the three scenarios. In all the cases, the resonantly enhanced final asymmetry is higher than the observed asymmetry of the Universe for $M_{1}\gtrsim 100$ GeV as discussed in the previous sections. In brief, the three scenarios are,
@@
    Case (a): Instantaneous RHN decay - Portrayed in Fig. 3, we classify this scenario by requiring $\Gamma_{N}\gg\Gamma_{\phi}$ which implies $T_{\rm N_{1}}\gg T_{\phi}$. This means that the RHNs are mostly produced at temperatures around $T_{\phi}$, when the scalar completely decays away and radiation starts dominating again. Those RHNs instantaneously decay to the SM bath. Even in the strong washout case $K\gg 1$, RHNs do not have time to thermalize with the radiation bath, thus the RHN decay process is strongly out-of-equilibrium. The evolution of the energy densities of different matter and radiation are shown in Fig. 3 for both weak ($K\ll 1$) and strong ($K\gg 1$) washout regime. We can see that for strong washout, the RHN density is lower than the weak washout case due to higher decay rate. From GW perspective, this case is only dependent of the total decay width of $\phi$ and independent of partial decay widths to RHNs and radiation, i.e. independent of the coupling $y_{R}$ given in Eq. (19). However the amount of leptogenesis depends on $y_{R}$.
@@
    Case (b): RHN Domination - For this scenario we need $y_{R}\sim 0$, and $\Gamma_{N}\ll\Gamma_{\phi}$ implying $T_{N_{1}}\ll T_{\phi}$ which means the RHN decay is delayed after production. As shown in Fig. 4, RHNs start to dominate the energy density as soon as the scalar decays away, with no radiation domination in between. When the RHNs decay out-of-equilibrium, radiation domination starts. If $K\ll 1$ the RHNs never thermalize with the radiation bath as seen in Fig. 4 (Left). If $K\gtrsim 1$, the RHNs eventually thermalize with the radiation bath due to their strong Yukawa interactions with the SM particles as shown in Fig. 4 (Right).
@@
    Case (c): Thermal leptogenesis - For large enough $y_{R}$ and for $T_{N_{1}}\ll T_{\phi}$, the RHNs produced from $\phi$ can quickly thermalize with the radiation bath and attain thermal abundance via 2-to-2 scattering process $N_{i}N_{i}\leftrightarrow\phi\leftrightarrow\overline{f}f$ (see Appendix B for details). This scenario is similar to thermal leptogenesis. The scalar will decay to both RHNs and radiation that will usher in an era of radiation domination, with the RHNs contributing to the radiation. In general, RHNs will also be produced thermally from the radiation bath via the scattering process. After attaining thermal abundance, eventually the RHNs freeze-out when the scattering process decouples, and then decay at $T_{N_{1}}$. Depending on their decay widths, these RHNs may or may not dominate the energy density prior to its decay. We are interested in the scenario when they do dominate. The necessary condition for that is Giudice:1999fb ; Berbig:2023yyy ,
    $T_{N_{1}}<T_{\rm dom}^{N_{1}}\lesssim 2\%M_{1}$ (32)
    which is the fermionic version of Eq. (25). The scalar dynamics adds nothing new to the final lepton asymmetry compared to the traditional thermal leptogenesis scenario. On the other hand, entropy injection now takes place in two stages: first at around $T_{\phi}$ when $\phi$ decays and secondly at RHN decay temperature $T_{N_{1}}$. Hence this case corresponds to 2-step entropy injection contrary to the non-thermal case, where entropy is injected in 1 step. A requirement for two-step entropy injection is,
    $T_{\phi}\gg T_{\rm dom}^{N_{1}}$ (33)
    such that after their non-thermal production from $\phi$, RHNs get enough time to thermalize then freeze-out before they start to dominate.
Figure 3: The evolution of the energy densities of the scalar $\rho_{\phi}$, RHN $\rho_{N}=\rho_{N_{1}}+\rho_{N_{1}}$ and radiation, $\rho_{R}$ with respect to the radiation temperature. Here we have taken $M_{1}=10^{5},M_{\phi}=3M_{1},T_{\phi}=10^{3}$ GeV for weak washout $K=2\times 10^{-4}$ (Left) and strong washout $K=50$ (Right). Both correspond to Case (a) mentioned at the beginning of Sec. IV. We can see RHNs instantaneously decay after production. The bottom panels in each of the images show the amount of entropy injected, $S=sa^{3}$ where $s\propto T^{3}$ is the entropy density defined in Eq. (24) and $a$ is the scale factor.
Figure 4: Same as Fig. 3 but for RHN domination scenario, as described in Case (b) of Sec. IV. For weak washout, $K\ll 1$ (Left), RHNs are always out-of-equilibrium. For strong washout, $K\gg 1$ (Right), RHNs get thermalized later due to their strong Yukawa interactions. The final asymmetry can be calculated numerically by solving the Boltzmann equations. The bottom panels show the entropy injection.

### v2

@@ III.2 The Boltzmann Equations
The relevant Boltzmann equations for the evolution of energy densities of the majoron, RHNs, radiation, and BLB-L asymmetry are,
$\dot{\rho}_{\phi}$ $=-3\mathcal{H}\rho_{\phi}-\Gamma_{\phi}\,(\rho_{\phi}-\rho_{\phi}^{\rm eq})$ (43)
$\dot{\rho}_{N_{1}}$ $=-3\mathcal{H}\left(\rho_{N_{1}}+p_{N_{1}}\right)+\Gamma_{\phi\rightarrow N_{1}N_{1}}\,(\rho_{\phi}-\rho_{\phi}^{\rm eq})-\left(\Gamma_{N_{1}}+\Gamma_{\rm scatt}\right)(\rho_{N_{1}}-\rho_{N_{1}}^{eq})$
$\dot{n}_{B-L}$ $=-3\mathcal{H}n_{B-L}-\epsilon\,\Gamma_{N_{1}}(n_{N_{1}}-n_{N_{1}}^{eq})-\left(\Gamma_{ID}+\Gamma_{ID}^{\rm scatt}\right)\,n_{B-L}$
$\dot{\rho}_{R}$ $=-4\mathcal{H}\rho_{R}+\left(\Gamma_{N_{1}}+\Gamma_{\rm scatt}\right)(\rho_{N_{1}}-\rho_{N_{1}}^{eq})$
where $\mathcal{H}$ is the Hubble parameter in cosmic time and dot denotes the derivative with respect to cosmic time and $\rho^{eq}_{X}$ and $n_{X}^{eq}$ represents the equilibrium energy and number density of the species $X$ respectively. Here we consider the pressure of the RHN, $p_{N_{1}}=\rho_{N_{1}}/3$ for $T>T_{\rm NR}$ where $T_{\rm NR}$ is defined in Eq. (37), and $p_{N_{1}}=0$ for $T\leq T_{\rm NR}$ , signifying the relativistic and non-relativistic evolution of the RHNs respectively. We will ignore the inverse decay term $\Gamma_{\phi}\rho_{\phi}^{\rm eq}$ since majoron decay temperature $T_{\phi}$ is much less than the majoron mass $M_{\phi}$, hence $\phi$ never comes into thermal equilibrium after it starts dominating. $\Gamma_{ID}$ is the inverse decay rate. See Appendix B for details.
In the above equations, we have written terms for decays, inverse decays and scattering processes of the non-thermally produced RHNs. Scatterings involving the majoron $lH\leftrightarrow N\phi$ are thermally suppressed at $T\ll M_{\phi}$. For $T_{\rm RH}\gg M_{1}$, thermal $N_{1}$ production via $\Delta L=1$ scatterings (e.g. $N_{1}l\leftrightarrow qt$) is subdominant to non-thermal production from $\phi$ decay, since $\phi$ dominates the energy density prior to decay. Scatterings of non-thermal $N_{1}$ with the plasma, described by $\Gamma_{\rm scatt}=2\Gamma_{Ss}+4\Gamma_{St}$ and the associated washout $\Gamma_{ID}^{\rm scatt}=\Gamma_{Ss}\frac{\rho_{N_{1}}}{\rho_{N_{1}}^{eq}}+2\Gamma_{St}$ Buchmuller:2004nz ; Zhang:2023oyo , become relevant only when $\rho_{N_{1}}\sim\rho_{R}$, which happens at temperatures around $T_{\phi}$ and $T_{N_{1}}$. Here $\Gamma_{Ss}$ and $\Gamma_{St}$ are the $s$- and $t$-shannel scattering rates involving top quark and gauge bosons. We solved the above Boltzmann equations with and without the scattering terms and found that, within the parameter space of interest, the inclusion of these effects modifies the final baryon asymmetry mostly by $10\%$. Only at $K\sim 20$, the change is $\sim 50\%$. This means $N_{1}$ decays and inverse decays dominate the dynamics compared to scatterings. A full kinetic treatment (see e.g. Hahn-Woernle:2009jyb ; Baker:2024udu ) is beyond the scope of this paper, hence we neglect scatterings in the rest of the analysis.
We solve the Boltzmann equations from the time when the majoron energy density overtakes the radiation energy density. Thus the initial energy density of the majoron is equal to the radiation density Kolb:1990vq ,
@@
We take $T_{\rm dom}$ from Eq. (34). The initial density of RHN is taken to be free parameter smaller than $\rho_{\phi}(T_{\rm dom})$ while initial BLB-L asymmetry is taken to be negligible. Further details on the numerical solution of the Boltzmann equations and calculation of the final $\kappa_{f}$ value are given in Appendix B. In Fig. 2, along with the thermal case, we show the enhancement in $\kappa_{f}$ in the non-thermal leptogenesis scenario by solving the Boltzmann equations. The solid magenta line and dashed green line show the enhancement of the final efficiency factor as a function of the washout parameter for $M_{1}=10^{5}$ GeV, $M_{\phi}=10^{7}$ GeV and $T_{\phi}=10^{5},3\times 10^{4}$ GeV in the non-thermal scenario. One can compare with thermal leptogenesis scenario (solid red and dashed blue lines). We notice that the value of washout parameter $K$ at which $\kappa_{f}$ is maximum increases with decreasing $T_{\phi}$ for the given parameters in Fig. 2, in contrast to results found in Hahn-Woernle:2008tsk ; Zhang:2023oyo . This is because, unlike Hahn-Woernle:2008tsk ; Zhang:2023oyo , we have considered the relativistic phase of RHNs in the Boltzmann equations and our RHNs decay width given in Eq. (39) is not thermally averaged.
At this point, we can numerically calculate the final baryon asymmetry in our model. However, we are not interested in the exact amount of asymmetry generated, since our goal is studying the GW probe of leptogenesis. In the next section, we discuss the classification of scenarios based on numerical results and analytical approximations, relevant to GW sensitivity.
Figure 2: Final efficiency factor $\kappa_{f}$ as a function of washout parameter $K$ in thermal and non-thermal leptogenesis. The solid blue and the dashed red curves represent a thermal and vanishing initial abundance of $N_{1}$ respectively in thermal leptogenesis. Here we used Eq. (16) and (17) for the blue curve while for the red curve, we use Eq. (15) for $K\ll 1$ and interpolate it to the thermal case in the $K\gg 1$ regime. The solid magenta and dashed green lines are solutions to the Bolzmann equations (43) for non-thermal leptogenesis with $M_{1}=10^{5}$ GeV, $M_{\phi}=10^{7}$ GeV for $T_{\phi}=10^{5}$ GeV, and $3\times 10^{4}$ GeV respectively.
@@ IV Classification of Scenarios
In this section, we identify three general non-thermal scenarios relevant for GW observations along the lines of Zhang:2023oyo . We present analytical estimates of the final asymmetry, equilibrium temperature when matter domination ends, and the dilution factor from entropy injection for each of the four scenarios. In brief, the four scenarios are,
@@
    Case (a): Instantaneous RHN decay - Portrayed in Fig. 3, we classify this scenario by requiring $\Gamma_{N}\gg\Gamma_{\phi}$ which implies $T_{\rm N_{1}}\sqrt{\frac{2M_{1}}{M_{\phi}}}\gg T_{\phi}$ using Eq. 39. This means that the RHNs are mostly produced at temperatures around $T_{\phi}$ when the majoron completely decays away and radiation starts dominating again. Those RHNs instantaneously decay to the SM bath because the lifetime of the highly relativistic RHNs, even after time dilation by Lorentz factor $\gamma_{N}=M_{\phi}/2M_{1}$, is still shorter than the lifetime of majoron $\phi$. In the non-relativistic RHN limit, i.e. $\gamma_{N}\rightarrow 1$, the condition for instantaneous RHN decay becomes $T_{N_{1}}\gg T_{\phi}$. In both relativistic and non-relativistic scenarios, RHNs do not have time to thermalize with the radiation bath even in the strong washout regime $K\gg 1$, thus the RHN decay process is strongly out-of-equilibrium. The evolution of the energy densities of different matter and radiation are shown in Fig. 3 for both weak ($K\ll 1$) and strong ($K\gg 1$) washout regimes. We can see that for a strong washout, the RHN density is lower than the weak washout case due to a higher decay rate of the RHNs. The final asymmetry in both the weak and strong washout regions can be estimated analytically by finding the value of $n_{N}$ at temperature $T_{\phi}$ by equating $\rho_{\phi}=\rho_{R}$ and $n_{N}=2\,n_{\phi}$ as Zhang:2023oyo ,
    $Y_{B}\equiv\frac{n_{B}-n_{\overline{B}}}{s}=\frac{3}{2}a_{\rm sph}\epsilon\frac{T_{\phi}}{M_{\phi}}$ (45)
    In this case, one can assume entropy is injected directly into the SM thermal bath at the time of majoron decay since RHN decay is instantaneous.
    Figure 3: The evolution of the energy densities of the majoron $\rho_{\phi}$, RHN $\rho_{N}=\rho_{N_{1}}$ and thermal radiation, $\rho_{R}$ with respect to the thermal radiation temperature. Here we have taken $M_{1}=10^{5},M_{\phi}=10^{7},T_{\phi}=10^{3}$ GeV for weak washout $K=10^{-2}$ (Left) and strong washout $K=10^{2}$ (Right). Both correspond to Case (a) mentioned at the beginning of Sec. IV. We can see RHNs instantaneously decay after production. The bottom panels in each of the images show the total entropy of the Universe, $S=s_{\rm total}a^{3}$ where $s_{\rm total}=s+s_{\rm N}$, with $s$ corresponding to the thermal sector entropy density, defined in Eq. (33) and $s_{N}$ being the entropy density of the decoupled RHN sector. Here $a$ is the scale factor.
@@
    Case (b): RHN Radiation Domination - For this scenario we need $\Gamma_{N}\ll\Gamma_{\phi}$ implying $T_{N_{1}}\sqrt{\frac{2M_{1}}{M_{\phi}}}\ll T_{\phi}$ which means the RHN decay is delayed after production. The RHNs dominate the energy budget as radiation after the majoron $\phi$ decays. The RHNs will decay relativistically at temperature $T_{N}^{\rm rel}$ when their decay rate becomes equal to the Hubble rate, i.e. $\Gamma_{N}(T_{N}^{\rm rel})\sim\mathcal{H}(T_{N}^{\rm rel})$. Considering $\mathcal{H}(T_{N}^{\rm rel})\sim(T_{N}^{\rm rel})^{2}/M_{\rm pl}$ and using Eq.s (36), (39) and (40), we obtain the decay temperature of RHNs to be,
    $T_{N}^{\rm rel}\sim\left(T_{N_{1}}^{2}\,T_{\rm NR}\right)^{1/3}$ (46)
    The necessary condition for RHN radiation domination is $T_{\rm NR}<T_{N_{1}}$ which also implies $T_{N}^{\rm rel}<T_{N_{1}}$ as shown in Fig. 4. The (Left) plot shows the RHN radiation domination for $K\ll 1$ whereas the (Right) shows RHN radiation domination for $K\gg 1$. In the weak washout regime ($K\ll 1$), one can calculate $n_{N}$ at $T_{N}^{\rm rel}$ by considering $\rho_{N}=\rho_{R}$, where $\rho_{N}$ can be written as $E_{N}^{\rm rel}n_{N}$. Using Eq. (41), we find that the final asymmetry turns out to be the same as Case (a) above, given in Eq. (45). For $K\gg 1$, the final asymmetry can be calculated by solving the Boltzmann equations. Since the RHNs decay relativistically, there is negligible entropy injection at $T_{N}^{\rm rel}$. However, significant entropy is injected into the RHN sector from the majoron decay at around temperature $T_{\phi}$.
    Figure 4: Same as Fig. 3 but for RHN radiation domination scenario, as described in Case (b) of Sec. IV, with $M_{\phi}=10^{7}$ and $T_{\phi}=10^{3}$ GeV assuming $M_{1}=10^{5}$ GeV, $K=10^{-5}$ (Left) and $M_{1}=10^{3}$ GeV, $K=100$ (Right). RHNs dominate as radiation starting from temperature $T_{\phi}$ till their decay at $T_{N}^{\rm rel}$ defined in Eq. (46). For weak washout, $K\ll 1$ (Left), RHNs never themalizes however the figure on (Right) shows RHN radiation domination for $K\gg 1$, where RHNs can thermalize later due to strong Yukawa couplings. The bottom panels show the entropy injection.
@@
    Case (c): RHN Matter Domination - We again consider $\Gamma_{N}\ll\Gamma_{\phi}$ for this scenario as Case (b), i.e. $T_{N_{1}}\sqrt{\frac{2M_{1}}{M_{\phi}}}\ll T_{\phi}$. However, unlike Case (b), here we require $T_{N_{1}}\ll T_{\phi}$, more specifically we are interested in the scenario where $T_{N_{1}}<T_{\rm NR}<T_{\phi}$ which translates to $\sqrt{K}<2T_{\phi}/M_{\phi}$. Thus strong washout $K>1$ is not possible in this scenario in our model. Relativistic RHNs first dominate the energy budget of the Universe as radiation after the decay of the majoron $\phi$. Then the RHNs become non-relativistic at $T_{\rm NR}$ and dominate as matter until their decay at temperature $T_{N_{1}}$. Hence there are two matter-dominated epochs separated by a radiation-dominated epoch and hence two stages of entropy injection in this case, first when the majoron decays to relativistic RHNs at $T_{\phi}$ and second when the RHNs decay at $T_{N_{1}}$. In the first stage, entropy is injected into the decoupled RHN sector and not into the SM thermal bath. When the RHNs decay out-of-equilibrium, the final radiation domination starts. If $K\ll 1$ the RHNs never thermalize with the radiation bath as seen in Fig. 5 (Left). The final BAU for $K\ll 1$ is estimated as Zhang:2023oyo ,
    $Y_{B}=\frac{3}{4}a_{\rm sph}\epsilon\frac{T_{N_{1}}}{M_{1}}\simeq 0.26\sqrt{K}\epsilon$ (47)
    In Fig. 6, we show the parameter space of the lightest RHN, spanned by $M_{1}$ and $K$ and the regions of successful non-thermal leptogenesis within it, i.e. $\eta_{B}\geq\eta_{B}^{\rm CMB}$, for different values of $M_{\phi}$ and $T_{\phi}$. The darker gray region at the bottom left is where sphaleron processes are turned off and leptogenesis is not efficient. The different colored regions correspond to different values of $M_{\phi}$ and $T_{\phi}$. For this plot we utilize analytical expressions given in Eq. (45), (47) for Case (a), Case (c), and Case (b) $K<1$ regions, and numerical solutions of the Boltzmann equations for Case (b) $K>1$ region.
    Figure 5: Same as Fig. 3 but for RHN matter domination scenario, described in Case (c) of Sec. IV, with $K=10^{-6}$, $M_{1}=10^{5}$ GeV, $M_{\phi}=10^{7}$ GeV and $T_{\phi}=10^{4}$ GeV. RHNs dominate as radiation first then as matter after temperature $T_{\rm NR}\sim 200$ GeV, defined in Eq. (37). The bottom panels show the two stages of entropy injection.
    Figure 6: Shaded regions of different colors correspond to successful non-thermal leptogenesis, $\eta_{B}\gtrsim\eta_{B}^{\rm CMB}$, in the parameter space of the lightest RHN for given values of $M_{\phi}$ and $T_{\phi}$. The dark gray region at the left bottom is where the RHN decay temperature is less than 100 GeV, hence sphaleron processes are suppressed and leptogenesis is not efficient.
Figure 7: Schematics of key events and temperatures for Case (a), (b) and (c). Inflation ends at reheating temperature $T_{\rm RH}$ after which $\phi$ remains in thermal equilibrium till its freeze out temperature $T_{f}^{\phi}$. Then $\phi$ becomes non-relativistic and start the first matter-domination epoch at $T_{\rm dom}\sim 1\%M_{\phi}$. This matter-domination ends when $\phi$ decays at $T_{\phi}$. For case (a), RHNs also instantaneuously decay at $T_{\phi}$ and leptogenesis occurs before the BBN temperature $T_{\rm BBN}$. In case (b), RHNs dominate as radiation from $T_{\phi}$ till their decay at $T_{N}^{\rm rel}$. In case (c), RHNs dominate as radiation till $T_{\rm NR}$ at which point they become non-relativistic and start the second MD epoch. This epoch ends at RHN decay temperature $T_{N_{1}}$ when leptogenesis occurs.